# SpikeFit: Towards Optimal Deployment of Spiking Networks on Neuromorphic Hardware

### v1

@@ 1 Introduction
The key characteristics of SpikeFit are: a) The values in the codebook (the set of $M$ discrete levels) are not fixed but shift to positions that are most advantageous for minimizing the global loss function b) The network’s underlying latent weights begin to cluster around these learned codebook values as training progresses c) Classic quantization scheme is applied to codebook values to reduce numerical precision d) Optimal clustering technique (CAT) allows to reduce performance loss on post-hoc weight quantization step acting like a quantization regularization technique (making the network less sensetive to noise or drift) e) The method allows to represent synaptic weights in accordance to available number of unique synaptic states (e.g. 4 available synaptic weight states in IBM TrueNorth [37]) while maintaining a low numerical precision of weight values d) The Fisher Spike Contribution (FSC) Pruning method precisely removes less important parts of the spiking network according to the Fisher Information of both its spiking activity, and its effect on a global loss function.
@@ 3 Quantized and Clustered SNN Baseline
In both baselines as well as any other SNN architectures that will be mentioned in this paper we implement the widely used The Leaky Integrate-and-Fire (LIF) neuron model to build a spiking network. The LIF neuron model is defined as follows:
The foundational continuous-time LIF model describes the membrane potential $U_{\mathrm{mem}}(t)$ of an RC circuit with the differential equation:
$\tau\frac{\mathrm{d}U_{\mathrm{mem}}(t)}{\mathrm{d}t}=-U_{\mathrm{mem}}(t)+R\,I_{\mathrm{in}}(t),$ (1)
where $\tau=RC$ is the membrane time constant. A forward Euler discretisation of (1) with a time step of $\Delta t$ yields the recurrence:
$U[t+1]=U[t]+\frac{\Delta t}{\tau}\Bigl(-U[t]+R\,I_{\mathrm{in}}[t]\Bigr).$ (2)
For deep learning applications, this is commonly simplified. Assuming $\Delta t=1$, $R=1$, and absorbing model parameters into learnable synaptic weights $W$, we arrive at the discrete LIF recurrence. The LIF neuron used in this work implements a soft-reset mechanism, which is particularly effective for surrogate gradient-based training. Its dynamics are defined as:
$U[t+1]=\beta\,U[t]+W\,X[t+1]-\beta\,S[t]\,U_{\mathrm{thr}},$ (3)
where $X[t]$ is the presynaptic spike train, $\beta=e^{-\Delta t/\tau}$ is the decay rate, and $S[t]=\mathbf{1}\!\bigl(U[t]>U_{\mathrm{thr}}\bigr)$ is the output spike [41]. This soft reset, scaled by $\beta$, contrasts with a hard reset ($S[t]\,U_{\mathrm{thr}}$), retaining essential spike timing dynamics while improving training stability.
@@
Clustered baseline method uses Artificial Neural Network (ANN) to Spiking Neural Network (SNN) conversion method to receive Spiking Neural Network architecture ready for full deployment on most neuromorphic hardware.
@@ 4.2 Fisher Spike Contribution (FSC) Pruning
Temporal and spiking specifics.
Because $y$ and $\delta$ are time-resolved, FSC integrates temporal coding by summing over $t$ in Eq. 9. In SNNs trained with surrogate gradients [17], $\delta$ captures the task-aligned sensitivity through time, making FSC a natural metric for identifying channels whose spikes causally influence the loss.
Computational cost.
FSC adds one backward pass over the limited saliency mini-batches. Its per-layer overhead is linear in $NTHW$ for the reductions and requires no SVD or eigendecomposition.
@@
Method Network Pruning Size (MB) Float Params Acc (%) Prec (%) Rec (%) F1 (%)
@@
Empirically, FSC consistently outperforms activity- or rank-based criteria (e.g., SVS) across sparsity regimes in our QP-SNN setting. We hypothesize three reasons: (i) Loss alignment: FSC is directly proportional to (a diagonal approximation of) the Fisher information $F_{cc}$ of a channel gate, hence it targets channels whose removal minimally increases expected loss, unlike SVS which is label-agnostic and rather measures representational complexity, but not task relevance. Two channels with similar SVS score can have vastly different impacts on the loss: one may spike during decisive moments (large $|\delta|$), while another spikes in background regions (small $|\delta|$). FSC distinguishes them by weighting spikes with instantaneous loss sensitivity. (ii) Temporal credit assignment: because $S_{c}^{\text{FSC}}$ integrates over $t$, it preserves channels that spike at behaviorally critical times (large $|\delta|$), a key aspect of SNN coding that static spatial rank cannot capture. (iii) Robustness to sign/quantization noise: squaring $\delta$ and $y$ avoids cancellations endemic to first-order Taylor and dampens noise introduced by quantization or surrogate-gradient stochasticity.
@@ 5 Experiment
Evaluation pipeline
Spiking Neural Networks often used to solve a set of problems in resource-constrained scenarios. Many of the problems could be decomposed into binary or multi-label classification of visual signals. For this reason, we decided to evaluate different SNN compression methods with multi-label classification task implementing widely known architectures in computer vision field (VGG-16, ResNet18) converted to their SNN versions by. We used standart benchmarks (datasets) in machine learning community for evaluation: CIFAR-10 (contains 60,000 32x32 color images in 10 different classes), ImageNet (contains over 14 million high-resolution images in 22,000 categories) and Imagenetta (a subset of ImageNet with 10 classes only).
@@
There are many efficient SNN compression methods introduced by other authors allowing to successfully deploy spiking models on specialized hardware. However, many of the works lack clear and objective evaluation methods as most experiments include task performance metrics only (accuracy, F1, and similar) showing whether the model handles the task well, but not representing the degree of improvement of the SNN model’s performance on neuromorphic hardware (or reflect metrics for evaluating network performance on the processor selectively, which makes evaluation difficult for other researchers). To fully estimate a success of both compression and deployment of SNN models researchers would require a specialized metric assessing complete contribution of compression methods to efficient deployment across experiments.
@@
Effectiveness of SNNs deployment on neuromorphic processors is characterized by the following variables: latency, energy use, model size, task performance. Thus, we decided to propose a simple DeployRatio (DR) metric to evaluate the effectiveness of methods mentioned in our work and include it in further experiments. We propose DeployRatio (DR), a composite metric to quantify the deployability of Spiking Neural Networks (SNNs) by jointly capturing task performance, latency, energy, and model footprint:
$\mathrm{DR}_{x}\;=\;\frac{\mathrm{Perf}_{x}}{L\times E\times S}\quad\text{with}\quad x\in\{\text{acc},\text{f1}\},$
where
$\mathrm{Perf}_{\text{acc}}=\frac{\mathrm{Accuracy}~[\%]}{100},\qquad\mathrm{Perf}_{\text{f1}}=\frac{\mathrm{F1}~[\%]}{100}.$
Here $L$ is inference latency (seconds per input), $E$ is energy per inference (mJ), and $S$ is model size (MB). Higher DR indicates a model that is simultaneously accurate, fast, energy efficient, and compact.
The components of the DeployRatio are:
  1. 1.
    Task performance $\mathrm{Perf}_{x}\in[0,1]$: either accuracy or macro-$\mathrm{F1}$ normalized to $[0,1]$ by dividing percentages by $100$.
  2. 2.
    Latency $L$ (s): wall-clock inference time averaged over multiple inputs/batches; we report mean latency over $B$ batches:
    $L\;=\;\frac{1}{B}\sum_{b=1}^{B}\Delta t_{b}.$
  3. 3.
    Model size $S$ (MB): computed from parameter count $P$ and effective precision $b$ (bits/weight):
    $S\;=\;\frac{P\cdot b}{8\cdot 10^{6}}~\mathrm{MB}.$
  4. 4.
    Energy $E$ (mJ): estimated from operation counts multiplied by per-operation energy constants (pJ/op) and converted to mJ.
DeployRatio metric is characterized by the following:
  1. 1.
    DR is dimensionless and monotonic: it increases with higher performance and decreases with greater latency, energy, or size.
  2. 2.
    $\mathrm{DR}_{\text{acc}}$ and $\mathrm{DR}_{\text{f1}}$ reflect differing task objectives (balanced vs. imbalanced classes).
@@
Dataset Network Method Precision Size (MB) Timestep Accuracy (%) F1 (%) DR (%) CIFAR-10 VGG-16 QP-SNN 8 $14.72$ 4 $86.58$ $86.91$ SpikeFit (CAT) 8 $14.72$ 4 89.14 89.99
@@ 5.1 Analysis of Results
The evaluation results across all experiments (Table 2) prove a consistent and compelling result on SpikeFit-optimized architectures performing best showing state-of-the-art performance across all tasks (ImageNet and CIFAR-10 multi-label classification). Moreover, SpikeFit allows for the most complete deployment of the SNN models by considering the limitation on unique synaptic states number required by some neuromorphic processors that focus on extreme energy efficiency (e.g. IBM TrueNorth) and showing the best performance according to the proposed DeployRatio (DR) metric.
@@ 6 Conclusion
SpikeFit training methodology demonstrated that combining a novel Clusterization-Aware-Training (CAT) method with quantization and the Fischer Spike Contribution (FSC) structured pruning scheme might be an effective way to deploy hardware-friendly SNNs that operate fast, with low-energy use at state-of-the-art performance. Hopefully, our contribution will motivate researchers to introduce methods offering even better energy-efficiency and performance on a wide range of neuromoprhic processors.

### v2

# RethinkingAI

@@ 1 Introduction
The key characteristics of SpikeFit are: a) The values in the codebook (the set of $M$ discrete levels) are not fixed but shift to positions that are most advantageous for minimizing the global loss function b) The network’s underlying latent weights begin to cluster around these learned codebook values as training progresses c) Optimal clustering technique (CAT) allows to reduce performance loss on post-hoc weight quantization step acting like a quantization regularization technique (making the network less sensetive to noise or drift) d) The method allows to represent synaptic weights in accordance to available number of unique synaptic states (e.g. 4 available synaptic weight states in IBM TrueNorth [37]) d) The Fisher Spike Contribution (FSC) Pruning method precisely removes less important parts of the spiking network according to the Fisher Information of both its spiking activity, and its effect on a global loss function.
@@ 3 Quantized and Clustered SNN Baseline
We employ the widely used Leaky Integrate-and-Fire (LIF) neuron model across all SNN architectures. The continuous-time LIF dynamics describe the membrane potential $U_{\mathrm{mem}}(t)$ of an RC circuit:
$\tau\frac{\mathrm{d}U_{\mathrm{mem}}(t)}{\mathrm{d}t}=-U_{\mathrm{mem}}(t)+R\,I_{\mathrm{in}}(t),$
where $\tau=RC$ is the membrane time constant. Using forward Euler discretization with step $\Delta t$:
$U[t+1]=U[t]+\tfrac{\Delta t}{\tau}\!\left(-U[t]+R\,I_{\mathrm{in}}[t]\right).$
For deep learning, we set $\Delta t=R=1$ and absorb parameters into learnable weights $W$, yielding a discrete form with soft reset:
$U[t+1]=\beta U[t]+WX[t+1]-\beta S[t]U_{\mathrm{thr}},$
where $X[t]$ is the presynaptic spike train, $\beta=e^{-\Delta t/\tau}$ the decay rate, and $S[t]=\mathbf{1}(U[t]>U_{\mathrm{thr}})$ the output spike [41]. The soft reset, scaled by $\beta$, preserves spike timing while enhancing training stability compared to a hard reset.
@@ 4.2 Fisher Spike Contribution (FSC) Pruning
Method Network Pruning Size (MB) Energy (mJ) Acc (%) Prec (%) Rec (%) F1 (%)
[…]
@@
Empirically, FSC consistently outperforms activity- or rank-based criteria (e.g., SVS) across sparsity regimes in our QP-SNN setting. We hypothesize the reason is the loss alignment: FSC is directly proportional to (a diagonal approximation of) the Fisher information $F_{cc}$ of a channel gate, hence it targets channels whose removal minimally increases expected loss, unlike SVS which is label-agnostic and rather measures representational complexity, but not task relevance. Two channels with similar SVS score can have vastly different impacts on the loss: one may spike during decisive moments (large $|\delta|$), while another spikes in background regions (small $|\delta|$).
@@ 5 Experiment
There are many efficient SNN compression methods introduced by other authors allowing to successfully deploy spiking models on specialized hardware. However, many of the works lack clear and objective evaluation methods as most experiments include task performance metrics only (accuracy, F1, and similar) showing whether the model handles the task well, but not representing the degree of improvement of the SNN model’s performance on neuromorphic hardware. To estimate a success of both compression and deployment of SNN models we propose a novel metric and include it in our experiments.
@@
We propose DeployRatio (DR), a simple composite metric to quantify the deployability of Spiking Neural Networks (SNNs) by jointly capturing task performance, latency, energy, and model footprint:
$\mathrm{DR}_{x}=\frac{\mathrm{Perf}_{x}}{L\,E\,S},\quad x\in\{\text{acc},\text{f1}\},\quad\mathrm{Perf}_{\text{acc}}=\tfrac{\text{Accuracy}[\%]}{100},\;\mathrm{Perf}_{\text{f1}}=\tfrac{\text{F1}[\%]}{100}.$
Here $L$ is inference latency (s/input), $E$ energy per inference (mJ), and $S$ model size (MB). A higher DR indicates a model that is accurate, fast, energy-efficient, and compact. The components are:
  • Task performance: $\mathrm{Perf}_{x}\in[0,1]$ from accuracy or macro-F1 (%/100).
  • Latency: $L=\frac{1}{B}\sum_{b=1}^{B}\Delta t_{b}$, mean inference time over $B$ batches.
  • Model size: $S=\frac{P\,b}{8\!\times\!10^{6}}$ MB, with parameter count $P$ and precision $b$ (bits/weight).
  • Energy: $E$ (mJ) estimated from operation counts and per-operation energy constants.
DR is dimensionless and monotonic—higher with better performance and lower with increased latency, energy, or size. $\mathrm{DR}_{\text{acc}}$ and $\mathrm{DR}_{\text{f1}}$ target different objectives (balanced vs. imbalanced classes).
@@
[…]
@@ 5.1 Analysis of Results
The evaluation results across all experiments (Table 2) prove a consistent result on SpikeFit-optimized architecture (VGG-16) performing best and showing state-of-the-art performance in the widely-known benchmark of multi-label classification on CIFAR-10 dataset. Moreover, SpikeFit allows for the most complete deployment of the SNN models by considering the limitation on unique synaptic states number required by some neuromorphic processors that focus on extreme energy efficiency (e.g. IBM TrueNorth) and showing the best performance according to the proposed DeployRatio (DR) metric.
@@ 6 Conclusion
SpikeFit training methodology demonstrated that combining a novel Clusterization-Aware-Training (CAT) method with quantization and the Fischer Spike Contribution (FSC) structured pruning scheme might be an effective way to deploy hardware-friendly SNNs that operate fast, with low-energy use at state-of-the-art performance. Hopefully, our contribution will motivate researchers to introduce methods offering even better energy efficiency and performance on a wide range of neuromoprhic processors.